An earthquake generates both transverse (S) and longitudinal (P) waves in the earth with speeds 4.5kms−1 and 8.0kms−1 respectively. A seismograph records that the first P-wave arrives 3.5 minutes earlier than the fust S-wave, From the seismograph, the epicentre of the earthquake is located at a distance
1
Expert's answer
2019-08-23T11:08:52-0400

Let the epicentre of the earthquake is located at a distance x.

The difference between the time taken by the P waves and S waves to reach the focus is


"3.5\ min=3.5(60)=210\ s."

Thus,


"\frac{x}{4.5}-\frac{x}{8}=210"

"x=2200\ km"